in 4x-5y=-17 what does y =
Found 2 solutions by richwmiller, Greenfinch:
Answer by richwmiller(17219)

We can solve for y in terms of x but can't get a number for y.
subtract 4x from both sides
-5y=-17-4x
divide by -5
y=(-17-4x)/-5

Answer by Greenfinch(383)

4x - 5y = -17
5y -4x = 17
5y = 4x + 17
y = (4x +17)/5